Every continuous function is differentiable.


True or false with full explanation.

1
Expert's answer
2022-06-03T12:51:47-0400

False.

Counterexample

The function "f(x)=|x|" is continuous on "(-\infin, \infin)."


"\lim\limits_{\Delta x\to0^{-}}\dfrac{f(0+\Delta x)-f(0)}{\Delta x}=\lim\limits_{\Delta x\to0^{-}}\dfrac{-\Delta-0}{\Delta x}=-1"

"\lim\limits_{\Delta x\to0^{+}}\dfrac{f(0+\Delta x)-f(0)}{\Delta x}=\lim\limits_{\Delta x\to0^{+}}\dfrac{\Delta-0}{\Delta x}=1"

"\lim\limits_{\Delta x\to0^{-}}\dfrac{f(0+\Delta x)-f(0)}{\Delta x}=-1"

"\not=1=\lim\limits_{\Delta x\to0^{+}}\dfrac{f(0+\Delta x)-f(0)}{\Delta x}"

Therefore


"\lim\limits_{\Delta x\to0}\dfrac{f(0+\Delta x)-f(0)}{\Delta x}"

does not exist.

Therefore the function "f(x)=|x|" is not differentiable at "x=0."